OpenStudy (anonymous):

the first side of the triangle measures 6 in. less than the second side, the third side is 5 in. more than the first and the perimeter is 20in . how long is each side..... I have a hard time with word problems

OpenStudy (anonymous):

im totally lost

OpenStudy (anonymous):

x be 14

OpenStudy (anonymous):

I got what u did I forget to take from both sides.. I struggle at this so bad thank u

OpenStudy (anonymous):

i though one side had to be 5 in less than the first

OpenStudy (anonymous):

oh wow, I could of sworn it was 6 less. That means my answer is wrong

OpenStudy (anonymous):

one side is 6in less the other is 5 in more

OpenStudy (anonymous):

 I think I might have did it right this time. Sorry, but it was hard to visualize the way you wrote it

OpenStudy (anonymous):

i know that's how they gave it me

OpenStudy (anonymous):

You could have drawn it earlier, and said my answer was wrong :)

OpenStudy (anonymous):

i was not sure .. so now how do i do that equation

OpenStudy (anonymous):

x = y - 6 z = 5+x x+y+z = 20

OpenStudy (anonymous):

so how do i get each side length

OpenStudy (anonymous):

You solve the system using substitution and/or elimination.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thank u so much